Circular No. 7220

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cq IN UGC 11268 A. V. Filippenko, University of California at Berkeley, reports: “In-

spection of an uncalibrated CCD spectrogram (range 330–1000 nm) obtained on July 9 UT with the 3-m Shane reflector at Lick Observatory confirms that SN 1999cq (cf. IAUC 7209) is a supernova. Its type is uncertain, but perhaps Ib or Ic entering the nebular phase. The Ca II near-infrared triplet and O I 777-nm are seen in emission, while Na D exhibits a P-Cyg profile. Hα is weak, if present. The object has faded by at least 2 mag since its discovery two weeks ago.”

NOVA IN NGC 6822 Filippenko further reports that a CCD spectrum (obtained as above)

confirms that the new object in NGC 6822 (cf. IAUC 7208, 7211, 7212) is indeed a nova. Prominent hydrogen Balmer and Fe II emission lines are visible.

V382 VELORUM C. E. Woodward, University of Wyoming; D. H. Wooden, Ames Re-

search Center, NASA; and R. K. Pina and R. S. Fisher, University of Florida, Gainesville, write: “We have discovered a strong [Ne II] 12.81μm line in the spectra of V382 Vel obtained on July 5.01 UT at the National Optical Astronomy Observatory’s 4-m Blanco telescope (+ OSCIR) at Cerro Tololo. Detection of the neon coronal emission line in the midinfrared spectra classifies V382 Vel as an ONeMg nova (cf. Gehrz et al. 1995, A.J. 110, 325). Also present in the spectra are H I recombination emission lines (9-7) and (7-6). The observed line fluxes, relative to H I (7-6) at 8.7 × 10^{−19} W cm^{−2}, are H I (9-7)/H I (7-6) = 0.5 and [Ne II]/H I (7-6) = 3.4. The ejecta velocity inferred from the H I (7-6) gaussian FWHM is 3300 km/s, consistent with earlier velocity estimates of Shore et al. (IAUC 7192). Continued infrared spectroscopy is urged.”

NOVAE IN M31 Corrigendum. On IAUC 5814 and 5861, the two positions given actu-

ally refer to the nucleus of M31; the offsets from the nucleus are correct in each case.

1999 July 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green